[Brahms-l] Shift report 20040302 00:00-08:00

From: Apache <apache@bnl.gov>
Date: Tue Mar 02 2004 - 07:55:07 EST
------------------------------------------------------------
Date: Tue Mar  2 07:55:07 2004, Shift: 00:00-08:00
Supervisor: BB, with: 
Log book#15 pages 71 - 71

Shift Summary:
------------------------------------------------------------
1. Beam for only one hour. AGS kicker extraction problem.

2. with FS at 8 deg, 1/2B and MRS at 40 deg, 1/2B

   Run 1720 35k trigger 2, 98k trigger 3

3. No detector problems

Magnets raised to full field in preparation for next running, and to identify potential heating problems.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue Mar 2 07:55:31 2004

This archive was generated by hypermail 2.1.8 : Tue Mar 02 2004 - 07:55:53 EST